OpenStudy (anonymous):

Two triangular roofs are similar. The ratio of the corresponding sides of these roofs is 2:3. If the altitude of the smaller roof is 4.5 feet, find the corresponding altitude of the larger roof. Round to the nearest hundredth.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f1a1419e4b04992dd228036 Your answer is here :D

jimthompson5910 (jim_thompson5910):

(smaller side)/(longer side) = ratio as a fraction 4.5/x = 2/3 4.5*3 = x*2 13.5 = 2x 2x = 13.5 x = 13.5/2 x = 6.75 So the longer altitude is 6.75 ft